OpenStudy (anonymous):

The first step in the Ostwald process for producing nitric acid is as follows: 4NH3(g) + 5O2(g) → 4NO(g) + 6H2O(g) If the reaction of 0.880 mol of NH3 and 0.469 mol of O2 yields 0.290 mol NO, what is the percent yield of this reaction?

OpenStudy (anonymous):

Do they specify which reagent or product you should find the yield?

OpenStudy (anonymous):

No. What I posted was all that was given.

OpenStudy (anonymous):

Ok. As they say ''yields NO''. Then I will assume it is this product. Again, we work on find the limiting and excess reagents. The molar ratio of NH3 to O2 is 4:5 or 1:1.25. If you are given 0.88mol of NH3, you are expected to have \[\frac{ 0.88*1.25 }{ 1 }=1.1mol\] of O2, but you are given less, which is 0.469mol. Therefore O2 is the limiting reagent and NH3 is in excess. Hence we work out between the limiting and the NO. The molar ratio of O2 to NO is 5:4 or 1:0.8 If we have 0.469 mol of O2, we are expected to have\[\frac{ 0.469*0.8 }{ 1 }=0.3752mol\] of NO, but we are given with 0.29mol of it. Ultimately, to find percentage yield, we use\[\frac{ Obtained yield*100 }{ expected yield }=\frac{ 0.29*100 }{ 0.3752 }=77.29%\]. The percentage yield is then 77.29%

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Thanks for explaining that. I have another question that is similar to this one but they gave grams instead of moles.

OpenStudy (anonymous):

Are you able to help me with it?

OpenStudy (anonymous):

Well, as I explained before, instead of working with ratio of moles, work with ratio of masses. Try it out then show me your working.

OpenStudy (anonymous):

What is the percent yield of chromium when 12.7 g of Cr is produced from the reaction of 40.0 g of Cr2O3 with 8.00 g of Al according to the following balanced chemical equation? 2Al + Cr2O3 → Al2O3 + 2Cr

OpenStudy (anonymous):

That's the question by the way.

OpenStudy (anonymous):

Cr2O3 has a mass of 152g

OpenStudy (anonymous):

The 12.7g Cr is the actual

OpenStudy (anonymous):

First, basing from the reactants side, find the limiting and excess reagents.

OpenStudy (anonymous):

Do you find that by doing 40.0g Cr2O3 * (2 mol Al/1 mol Cr2O3)?

OpenStudy (anonymous):

No. Do not mix moles with masses. Follow my steps.

OpenStudy (anonymous):

First consider the Mr of the reactants.

OpenStudy (anonymous):

Al is 27g Cr2O3 is 152g, like you told me.

OpenStudy (anonymous):

According to the reaction, (2*27) or 54g of Al are to 152g of Cr2O3, right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

If you have 8g of Al, it is expected for it to react with \[\frac{ 8*152 }{ 54 }=22.5g \] of Cr2O3, but you are given 40.0g of it. Can you tell me now which is the limiting and which is in excess?

OpenStudy (anonymous):

It would be Al since there is only 22.5g of it compared to the 40.0g of Cr2O3

OpenStudy (anonymous):

Correct! So now we work with the limiting reagent ratio to that of Cr. Try this step out. I will be back!

OpenStudy (anonymous):

Well? Finding hard?

OpenStudy (anonymous):

Cr2O3 = 22.535 Cr = 15.418

OpenStudy (anonymous):

Or did I do that completely wrong?

OpenStudy (anonymous):

Because I did (8.00g Cr2O3)(mol Al)(1 mol Cr2O3)(152g Cr2O3)/(26.98g Al)(2 mol Al)( mol Cr2O3) = 22.535

OpenStudy (anonymous):

And (8.00g Al)(mol Al)(2 mol Cr)(52g)/(26.98g Al)(2 mol Al)( mol Cr) = 15.418

OpenStudy (anonymous):

Nice one!

OpenStudy (anonymous):

So that's all right?

OpenStudy (anonymous):

Finally use the formula of percentage yield that I gave previously and work out the percentage.

OpenStudy (anonymous):

Yes, you got it! Now towards the final step. Find the percentage.

OpenStudy (anonymous):

The 15.418 is the theo right?

OpenStudy (anonymous):

Since the problem is looking for Cr

OpenStudy (anonymous):

Yes. That will be the expected yield.

OpenStudy (anonymous):

So (12.7)/(15.418) *100 = 82.37%

OpenStudy (anonymous):

That's it. You got it.

OpenStudy (anonymous):

Thank you so much for your help! You're so awesome!

OpenStudy (anonymous):

My pleasure.